# Learning positional encodings in transformers depends on initialization

### v3

@@ A.1 Positional encoding definitions
Below, we provide the formal definitions for the common PEs that we evaluated learnable PEs against: 1d-fixed, 2d-fixed, relative, and rope.
@@
relative [Shaw et al., 2018]. relative PE modifies standard self-attention to incorporate the relative positions of tokens. This implies that calculation of PE is wrapped within the self-attention module. The relative position embedding parameter between a token at position $i$ and $j$ is $a_{j-1}$. In brief, self attention is then modified to include relative position information by modifying attention between tokens $i$ and $j$ as
@@
rope [Su et al., 2022]. rope applies a rotation to the token embeddings based on their positions in a higher dimensional space. For a token at position $p$ with an embedding $x$, let $x=[x_{1},x_{2},...,x_{d}]$, where $d$ is even. Then, for each pair of dimensions, apply the rotation
@@ A.3 Modularity and clustering in network data
Figure A10: Comparing top-performing learnable PE models (learn-0.2) with commonly-used PE schemes. A) We compare the generalization performance of learn-0.2 model with models using common PEs. Notably, we included a 2d-fixed PE as the “ground truth”, since it obeyed the 2D organization of the LST task. The learn-0.2 model outperformed all other PE models. B) We compared the attention maps of each model to those derived from the “ground truth” 2d-fixed model. C,D) We found that aside from the 1d-fixed model (which does not generalize well), learn-0.2 learned the closest attention map to the 2d-fixed model. This was expected, since the 2d-fixed and 1d-fixed PEs are highly similar by design (the baseline cosine similarity between the two schemes is 0.72). In contrast, the learn-0.2 learned an attention map that was highly similar to the 2d-fixed model, despite having no similarity to the 2d-fixed PE scheme at initialization (cosine at initialization = 0.00). (We also note that the rope model had high baseline similarity to the 2d-fixed, since by construction, a component of the rope encoding is highly similar to the 1d-fixed PE scheme.) Thus, we found that despite having no prior bias towards the 2d-fixed PE scheme, a small-norm initialized learnable PE is capable of learning an attention map that approximates an attention map derived from the ground truth PE.
Figure A11: A) We injected noise with different amplitudes into each token embedding, and assessed downstream generalization performance. B) Generalization performance with noisy inputs across all initializations for learnable PEs, and C) common PEs. D) Average performance across noisy inputs (collapsing across rows in B). E) Besides the ground truth model (2d-fixed), the learn-0.2 model was most robust to perturbation compared to common PEs.
@@
Figure A15: Model performances when incorporating multihead (2 and 4) attention mechanisms. Overall, we find that adding attention heads tends to reduce generalization performance across the board. The only architecturein which we see improvements are models with rope PE. Nevertheless, despite their improvements, rope models still do not outperform models with relative PEs, nor do they outperform learn-0.2 models with a single attention head (single head learn-0.2 performance = 95.6%; Table 1). We also notably see that learn-0.2 models tend to degrade in their performance, likely due to the increase in free parameters. A,B) Performance of models with 2 attention heads. C) Performance and of models with learnable PEs (2 attention heads). D) Training trajectories for example model architectures (2 attention heads). E,F) Performance of models with 4 attention heads. G) Performance of models with learnable PEs (4 attention heads). H) Training trajectories for example model architectures (4 attention heads).

### v4

@@ A.1 Positional encoding definitions
Below, we provide the formal definitions for the common PEs that we evaluated learnable PEs against: 1d-fixed, 2d-fixed, 1d-relative, and 1d-rope.
@@
1d-relative [Shaw et al., 2018]. 1d-relative PE modifies standard self-attention to incorporate the relative positions of tokens. This implies that calculation of PE is wrapped within the self-attention module. The relative position embedding parameter between a token at position $i$ and $j$ is $a_{j-1}$. In brief, self attention is then modified to include relative position information by modifying attention between tokens $i$ and $j$ as
@@
1d-rope [Su et al., 2022]. 1d-rope applies a rotation to the token embeddings based on their positions in a higher dimensional space. For a token at position $p$ with an embedding $x$, let $x=[x_{1},x_{2},...,x_{d}]$, where $d$ is even. Then, for each pair of dimensions, apply the rotation
@@ A.3 Modularity and clustering in network data
Figure A9: Comparing top-performing learnable PE models (learn-0.2) with commonly-used PE schemes. A) We compare the generalization performance of learn-0.2 model with models using common PEs. Notably, we included a 2d-fixed PE as the “ground truth”, since it obeyed the 2D organization of the LST task. The learn-0.2 model outperformed all other PE models. B) We compared the attention maps of each model to those derived from the “ground truth” 2d-fixed model. C,D) We found that aside from the 1d-fixed model (which does not generalize well), learn-0.2 learned the closest attention map to the 2d-fixed model. This was expected, since the 2d-fixed and 1d-fixed PEs are highly similar by design (the baseline cosine similarity between the two schemes is 0.72). In contrast, the learn-0.2 learned an attention map that was highly similar to the 2d-fixed model, despite having no similarity to the 2d-fixed PE scheme at initialization (cosine at initialization = 0.00). (We also note that the 1d-rope model had high baseline similarity to the 2d-fixed, since by construction, a component of the 1d-rope encoding is highly similar to the 1d-fixed PE scheme.) Thus, we found that despite having no prior bias towards the 2d-fixed PE scheme, a small-norm initialized learnable PE is capable of learning an attention map that approximates an attention map derived from the ground truth PE.
Figure A10: A) The brain is organized into functional networks that are spatially distributed throughout the brain in 3D space. B) We sought to understand the influence of PE on a model’s ability to predict contemporaneous brain activity. This can be formalized as a masked prediction task, where a transformer is trained to predict contemporaneous, spatially masked brain activity. To test for generalization, we evaluated the MSE on a validation dataset, which involved predicting the masked activity of data collected from different participants with increased masking (90%). (Models were trained with a 50% masking. See Figures A11,A12,A13 for results with 15%, 75%, and 90% masked training objectives.) C) We found that transformers endowed with different PE schemes converged to different MSE loss values. D) Notably, learnable PEs initialized with small norms (learn-0.1 and learn-0.2) achieved to the lowest MSE for both training and validation datasets. (X-axis is sorted from highest to lowest MSE.)
@@
Figure A14: Model performances when incorporating multihead (2 and 4) attention mechanisms. Overall, we find that adding attention heads tends to reduce generalization performance across the board. The only architecturein which we see improvements are models with 1d-rope PE. Nevertheless, despite their improvements, 1d-rope models still do not outperform models with 1d-relative PEs, nor do they outperform learn-0.2 models with a single attention head (single head learn-0.2 performance = 95.6%; Table 1). We also notably see that learn-0.2 models tend to degrade in their performance, likely due to the increase in free parameters. A,B) Performance of models with 2 attention heads. C) Performance and of models with learnable PEs (2 attention heads). D) Training trajectories for example model architectures (2 attention heads). E,F) Performance of models with 4 attention heads. G) Performance of models with learnable PEs (4 attention heads). H) Training trajectories for example model architectures (4 attention heads).
Figure A15: Control models. We assessed two alternative artificial neural network architectures: A vanilla multilayer perceptron (MLP), and a bidirectional Long Short-Term Memory (LSTM) model. Each model contained four layers and 160 hidden units each. Overall, we found that compared to the 2d-fixed PE transformer model, the MLP and LSTM model could not generalize the LST task.
Figure A16: Supplementary data associated with Fig. 4. We report the training and validation MSE evaluation at the end of training for the stochastic network simulation. A) Training converged rapidly across all models. B) Across all the models / PE variants, there was little difference in MSE across training and validation distributions for this task. This is because all models optimally converged. In addition, all models could not converge to near zero due to the stochastic nature of the simulation (which cannot be perfectly modeled).